a. Show = b. Show E{(X-EX)(Y-EY)}=E{(X-EX)Y}

a. Show B5dRk+6sL5nzdiZ2dQCFD+AHfqbUPGQrwiAAAAAE= yQpjfE9lAAAAAElFTkSuQmCC

b. Show E{(X-EX)(Y-EY)}=E{(X-EX)Y}
Answer #1

to show : Elnion 2 n ě ni Grineu i uns e hi-n) 2 Where u is the mean, ie a = ê ni - Now, oferining the brackets, & (ui-u)2 =& ni? + (ne) ž - 22 u ni SWI SW SWA į niz - u Žni = RHS Hently his =RHS heuce proved. to proue! E(X – E(x) XY-E(4)= E$(x = EDELIA the above 2 properties, we get > = E(xy) - Elv) EX2 - E(X) ELY) + EXEN) = E(XY) – E(X) ECV) E3C x= EW)) y 4 - RHS hea .